There is a 5% chance that the mean reading speed of a random samlple of 19 second grade studetns will exceed what value? u(x) =93 and SD = 2.294
Use technology to find the sample mean x for which there is an area of 0.05 to the right of the normal curve with u(x) of 93 and SD 2.294

Answer by stanbon(75887)

There is a 5% chance that the mean reading speed of a random sample of 19 second grade studetns will exceed what value? u(x) =93 and SD = 2.294
Use technology to find the sample mean x for which there is an area of 0.05 to the right of the normal curve with u(x) of 93 and SD 2.294
-----------
Find the z-value with a left tail of 95%
invNorm(0.95) = 1.645
------
Find the corresponding reading speed using x = z*(s/sqrt(n))+u
rs = 1.645*(2.294/sqrt(19) + 93 = 94.1
===============================
Cheers,
Stan H.
===============